OpenStudy (help_people):

Create a rational function with a linear binomial in both the numerator and denominator. Part 1. Graph your function using technology. Include the horizontal and vertical asymptotes and the x- and y-intercepts on your graph. Label the asymptotes and intercepts. Part 2. Show all work to identify the vertical asymptote, the x-intercepts, and the y-intercept.

OpenStudy (help_people):

@welshfella @LynFran

OpenStudy (help_people):

@adxpoi

OpenStudy (welshfella):

ax + b ------ cx - d plug in values for a b c and d ( not = 0)

OpenStudy (help_people):

a=1 b=2 c=3 d4? is that good? @welshfella srry i was inactive :(

OpenStudy (welshfella):

yes so we have x + 2 ----- 3x + 4 you can use desmos to graph this

OpenStudy (welshfella):

https://www.desmos.com/calculator/5qfhpf8bgt asymptotes are breaks in the graph

OpenStudy (welshfella):

you see that the x intercept is at (-2,0)

OpenStudy (help_people):

ok

OpenStudy (welshfella):

what is the y-intercept

OpenStudy (help_people):

0.5

OpenStudy (help_people):

x= -2

OpenStudy (welshfella):

y = 0.5 they want the coordinates of the point which is x = 0 y = 0.5 or (0,0.5)

OpenStudy (help_people):

ok

OpenStudy (welshfella):

y intercept is at ( 0 , 0.5)

OpenStudy (help_people):

ok

OpenStudy (welshfella):

now the vertical asymptote is the line between where the vertical parts of the go 'off the page'

OpenStudy (welshfella):

thats between x = -1 and x = -2 and its difficult to get from the graph vertical asymptotes occur when the denominator is zero so we can work it ouT:- 3x + 4 = 0 3x = -4 x = -4/3 or about -1.33

OpenStudy (help_people):

should i put x-intercept also ?

OpenStudy (welshfella):

yea we 've done that

OpenStudy (welshfella):

to work it out we put (x + 2) / (3x + 4) = 0 so x+ 2 = o and x = -2 x intercept at (-2 , 0)

OpenStudy (help_people):

how would you work out the intercept

OpenStudy (help_people):

*y-intercept

OpenStudy (welshfella):

put x = 0 in ( x +2) / (3x + 4)

OpenStudy (welshfella):

because at the intercepts x = 0

OpenStudy (help_people):

so (0+2) /3(0) + 4)? (2)/ (0+4) 2/4 1/2?

OpenStudy (welshfella):

yes - which is what we read off on the graph

OpenStudy (help_people):

and also the vertic asymptotes is -4/3 OR about -1.33 right but what about the horizontal ?

OpenStudy (welshfella):

i cant remember !! lol hold on

OpenStudy (help_people):

and ok :)

OpenStudy (help_people):

srry I'm throwing everything @ you!

OpenStudy (welshfella):

i gotta go right now as you see form the graph its about y = 0.6

OpenStudy (help_people):

really quick whats the horizontal retricem and show me plesase who to solve